# Testing by Betting while Borrowing and Bargaining

### v1

@@ Abstract
Testing by betting has been a cornerstone of the game-theoretic statistics literature. In this framework, a betting score (or more generally an e-process), as opposed to a traditional p-value, is used to quantify the evidence against a null hypothesis: the higher the betting score, the more money one has made betting against the null, and thus the larger the evidence that the null is false. A key ingredient assumed throughout past works is that one cannot bet more money than they currently have. In this paper, we ask what happens if the bettor is allowed to borrow money after going bankrupt, allowing further financial flexibility in this game of hypothesis testing. We propose various definitions of (adjusted) evidence relative to the wealth borrowed, indebted, and accumulated. We also ask what happens if the bettor can “bargain”, in order to obtain odds bettor than specified by the null hypothesis. The adjustment of wealth in order to serve as evidence appeals to the characterization of arbitrage, interest rates, and numéraire-adjusted pricing in this setting.
@@ 1 Introduction
In the standard setup of game-theoretic statistics, bets are offered to a statistician to test the verity of a null hypothesis [Shafer, 2021]. Before each outcome is revealed, the statistician places an arbitrary fraction of their wealth to bet on its value, at odds specified by the null hypothesis. The total wealth of the gambler can be interpreted as the amount of evidence available against the null.
In particular, we consider the mean estimation set-up studied by various authors including Shafer and Vovk [2005, Section 3] and Waudby-Smith and Ramdas [2024], where there is an infinite sequence of bounded random variables with a common conditional mean $\mu$. We are chiefly concerned with the special case where these random variables are coin tosses; that is, they are i.i.d. drawn from some distribution on {±$\{\pm 1\}$. The fairness of the coin, $H_{0}:\mu=1/2$, is to be tested. Extending our exposition to the general bounded case with a common conditional mean is straightforward.
@@
In this paper, however, we allow the betting statistician to borrow interest-free before placing each bet without being forced to quit the game in the possible event of indebtedness, and study if the emergence of evidence is changed in any way correspondingly. Further, we discuss the related scenario where the bets are underpriced such that arbitrage is possible without evidence.
Throughout the paper, we consider a fixed measurable space $(\Omega,\mathcal{A})$ with measurable functions $X_{1},X_{2},\dots:\Omega\to\{\pm 1\}$, construed as repeated coin tosses; and denote by $\{\mathcal{F}_{t}\}$ the canonical filtration they generate, i.e., $\mathcal{F}_{t}=\sigma(X_{1},\dots,X_{t})$. We denote by $\mathbb{P}_{p}$ the probability measure on $(\Omega,\mathcal{A})$ such that $X_{1},X_{2},\dots\overset{\mathrm{iid}}{\sim}\operatorname{Rad}_{p}$ (Rademacher ±$\pm 1$-valued random variables with probability of 1 being $p$), and $\mathbb{E}_{p}$ the corresponding expectation or conditional expectation.
2 Review of Testing by Betting (without Borrowing)
@@
The classical set-up of testing by betting pioneered by Shafer and Vovk [2005] involves the following game with two players, Casino [^1] This player is referred to as Forecaster in the original work, but in our work, Casino seems a better analogy. and Statistician. Casino tosses a coin repeatedly, realized as the random variables $X_{1},X_{2}\dots$ mentioned above, and Statistician does not believe that the coin is fair. At each time $t=1,2,\dots$, Casino offers a double-or-nothing bet to Statistician on the upcoming coin toss $X_{t}$. Statistician is allowed to place any fraction (between 0 and 1) of their wealth on either heads or tails. Formally speaking, let $\{\lambda_{t}\}_{t\geqslant 1}$ be a process bounded in $[-1,1]$, predictable with respect to the filtration $\{\mathcal{F}_{t}\}_{t\geqslant 0}$, where $\lambda_{t}$ the fraction of wealth Statistician bets on $X_{t}=1$. Let $W_{0}=1$ be the initial wealth of Statistician, whose wealth process therefore evolves as
$W_{t}=W_{t-1}\cdot(1+\lambda_{t}X_{t})$ (1)
over the course of the game.
@@
Under $\mathbb{P}_{1/2}$, i.e., when the coin is fair, $\mathbb{E}_{1/2}(X_{t})=0$, implying that the wealth process $\{W_{t}\}$ is a nonnegative martingale. Ville’s inequality [Ville, 1939] then implies that, for all $x>1$,
$\mathbb{P}_{1/2}\left(\sup_{t}W_{t}\geqslant x\right)\leqslant 1/x;$ (2)
@@
$\mathbb{P}_{1/2}\left(W_{\tau}\geqslant x\right)\leqslant 1/x.$ (3)
Along with the optional stopping theorem for nonnegative martingales, which states that $\mathbb{E}_{1/2}W_{\tau}\leqslant 1$ for any stopping time $\tau$, both inequalities above lend credibility to the interpretation that Statistician’s wealth $W_{t}$ can be seen as the evidence against the fairness of the coin $\mathbb{P}_{1/2}$. Inequality (2) states that if the coin was fair, it would be unlikely for Statistician ever to be wealthy, and thus a large value of $W_{t}$ at any time means that the coin is unlikely to be fair. Inequality (3) states that Statistician may employ any stopping rule and a large final stopped wealth discredits $\mathbb{P}_{1/2}$. Related to both is the concept of e-values, which are nonnegative random variables with expected values at most 1. In particular, $W_{\tau}$ is an e-value for any stopping time $\tau$.
@@
3 Evidence in Borrowed Betting
We now suppose before placing the bet on $X_{t}$ at time $t$, Statistician borrows [^2] It does not matter whom Statistician borrows the money from. $\beta_{t}$ amount of money, increasing (or decreasing, if $\beta_{t}<0$) their current (gross) wealth $W_{t-1}$ to $W_{t-1}+\beta_{t}$, of which Statistician then bets a $\lambda_{t}$ fraction on $X_{t}=1$. Formally, let $\{\beta_{t}\}_{t\geqslant 1}$ and $\{\lambda_{t}\}_{t\geqslant 1}$ be two predictable processes on $\{\mathcal{F}_{t}\}_{t\geqslant 0}$. We assume $\lambda_{t}\in[-1,1]$ and $\beta_{t}\geqslant-W_{t-1}$ for each $t$. In particular, we allow $\beta_{t}$ to be negative which could correspond to Statistician “paying back” earlier debts. Now, Statistician’s wealth process is defined as
$W_{0}=1,\quad W_{t}=(W_{t-1}+\beta_{t})\cdot(1+\lambda_{t}X_{t}),$ (4)
a nonnegative process adapted to $\{\mathcal{F}_{t}\}$.
To further clarify the set-up, we introduce the liabilities process denoting the total borrowed amount up to time $t$,
@@ 3 Evidence in Borrowed Betting
Recall that in the previous setting without borrowing in Section 2, the wealth process is a nonnegative martingale under $\mathbb{P}_{1/2}$. In our current case with borrowing, we have the following proposition on the more general relation between these three processes.
@@
The net wealth process $\{N_{t}\}$ is a martingale on $\{\mathcal{F}_{t}\}$ under $\mathbb{P}_{1/2}$. Further, $W_{t}=N_{t}+L_{t}$ is the Doob decomposition of the process $\{W_{t}\}$ into the martingale $\{N_{t}\}$ and the predictable process $\{L_{t}\}$. Consequently, $\{W_{t}\}$ is a submartingale if and only if the amount borrowed each time $\beta_{t}\geqslant 0$.
@@
$\mathbb{E}_{1/2}(N_{t}|\mathcal{F}_{t-1})=(W_{t-1}+\beta_{t})\cdot\mathbb{E}_{1/2}(1+\lambda_{t}X_{t}|\mathcal{F}_{t-1})-L_{t}=W_{t-1}+\beta_{t}-L_{t}=N_{t-1}$, therefore $\{N_{t}\}$ is a $\mathbb{P}_{1/2}$-martingale. It is easy to see that $\{L_{t}\}$ is predictable and $L_{0}=0$. Therefore, $W_{t}=N_{t}+L_{t}$ is the Doob decomposition of $\{W_{t}\}.$ The final claim is immediate. ∎
We would like to study how Statistician, now holding both wealth $W_{t}$ and liabilities $L_{t}$, can quantify the evidence against $\mathbb{P}_{1/2}$, based on either $W_{t}$ or $N_{t}$ or both. To generalize our previous discussion on interpreting evidence via inequalities (2) and (3), let us introduce the following definitions.
@@
Let $\tau$ be a stopping time on $\{\mathcal{F}_{t}\}$ and $a>0,b\geqslant 0$ be constants. We say that an $\mathcal{F}_{\tau}$-measurable random variable $E$ is an $(a,b)$-tail evidence against $\mathbb{P}_{1/2}$ at $\tau$ if the tail bound
$\mathbb{P}_{1/2}(E\geqslant x)\leqslant\frac{a}{x-b}$ (7)
@@
Random variables that are measurable with respect to the stopped $\sigma$-algebra $\mathcal{F}_{\tau}$ are often constructed by stopping at $\tau$ a process adapted to $\{\mathcal{F}_{t}\}$. This definition of tail evidence above follows the same reasoning that “if $\mathbb{P}_{1/2}$ holds, $E$ is unlikely to be large” as discussed in Section 2; the e-value $W_{\tau}$ back in (3), per Definition 3.2, is a $(1,0)$-tail evidence against $\mathbb{P}_{1/2}$ at $\tau$, and the constants $a,b$ in the definition are seen as “discount factors” of the evidence value $E$. More generally, if $a^{-1}(E-b)$ is an e-value, then $E$ is an $(a,b)$-tail evidence. Conversion from an $(a,b)$-tail evidence to a p-value for testing the null $\mathbb{P}_{1/2}$ can be done via $E\mapsto\frac{a}{(E-b)^{+}}$.
@@
$\mathbb{P}_{1/2}\left(\sup_{t}E_{t}\geqslant x\right)\leqslant\frac{a}{x-b},$ (8)
we say that it is a $(a,b)$-sequential tail evidence against $\mathbb{P}_{1/2}$.
For example, Ville’s inequality states that any nonnegative $\mathbb{P}_{1/2}$-supermartingale $\{M_{t}\}$ is an $(M_{0},0)$-sequential tail evidence. The stopping time-dependent Definition 3.2 can be converted to the time-maximal Definition 3.3, if the same constant pair $(a,b)$ works for all bounded stopping times, as shown in the following proposition.
@@
Let $\{E_{t}\}$ be a process adapted to $\{\mathcal{F}_{t}\}$. It is an $(a,b)$-sequential tail evidence against $\mathbb{P}_{1/2}$ if and only if for any bounded stopping time $\tau$, $E_{\tau}$ is an $(a,b)$-tail evidence against $\mathbb{P}_{1/2}$ at $\tau$.
@@
First, suppose $E_{\tau}$ is an $(a,b)$-tail evidence at $\tau$ for any bounded stopping time $\tau$. Let $\sigma$ be the stopping time $\inf\{t:E_{t}\geqslant x\}$, where by convention $\inf\varnothing=\infty$. Then, since $E_{\sigma\wedge T}$ is an $(a,b)$-tail evidence against $\mathbb{P}_{1/2}$ at $\sigma\wedge T$,
$\mathbb{P}_{1/2}\left(\sup_{t\leqslant T}E_{t}\geqslant x\right)=\mathbb{P}_{1/2}(E_{\sigma\wedge T}\geqslant x)\leqslant\frac{a}{x-b}.$ (9)
@@
$\mathbb{P}_{1/2}\left(\sup_{t}E_{t}\geqslant x\right)=\mathbb{P}_{1/2}\left(\bigcup_{T=1}^{\infty}\left\{\sup_{t\leqslant T}E_{t}\geqslant x\right\}\right)\leqslant\frac{a}{x-b},$ (10)
@@
Next, suppose $\{E_{t}\}$ is an $(a,b)$-sequential tail evidence and let $\tau\leqslant T$ be a bounded stopping time. Then $\mathbb{P}_{1/2}(E_{\tau}\geqslant x)\leqslant\mathbb{P}_{1/2}(\sup_{t\leqslant T}E_{t}\geqslant x)\leqslant\frac{a}{x-b}$, concluding that $E_{\tau}$ is an $(a,b)$-tail evidence at $\tau$. ∎
@@
The net wealth martingale $\{N_{t}\}$ satisfies the supermartingale optional stopping theorem with respect to stopping times $\tau$ and 0; that is, $\mathbb{E}_{1/2}N_{\tau}\leqslant N_{0}=1$.

### v2

@@ Abstract
Testing by betting has been a cornerstone of the game-theoretic statistics literature. In this framework, a betting score (or more generally an e-process), as opposed to a traditional p-value, is used to quantify the evidence against a null hypothesis: the higher the betting score, the more money one has made betting against the null, and thus the larger the evidence that the null is false. A key ingredient assumed throughout past works is that one cannot bet more money than they currently have. In this paper, we ask what happens if the bettor is allowed to borrow money after going bankrupt, allowing further financial flexibility in this game of hypothesis testing. We propose various definitions of (adjusted) evidence relative to the wealth borrowed, indebted, and accumulated. We also ask what happens if the bettor can “bargain”, in order to obtain odds better than specified by the null hypothesis. The adjustment of wealth in order to serve as evidence appeals to the characterization of arbitrage, interest rates, and numéraire-adjusted pricing in this setting.
@@ 1 Introduction
In the standard setup of game-theoretic statistics, bets are offered to a statistician to test the verity of a null hypothesis [Shafer, 2021]. Before each random outcome is revealed, the statistician places an arbitrary fraction of their wealth to bet on its value, at odds specified by the null hypothesis. The total wealth of the gambler can be interpreted as the amount of evidence available against the null.
To illustrate this concept, we may first consider the mean estimation set-up studied by various authors, including Shafer and Vovk [2005, Section 3] and Waudby-Smith and Ramdas [2024], where there is an infinite sequence of bounded random variables with a common conditional mean $\mu$. We assume these random variables are $\{-1,1\}$-valued, and that the null hypothesis is $H_{0}:\mu=0$. We shall write out the full setup in Section 2, and explain how the “wealth” in this setup admits a clear economic interpretation of betting on the outcome of coin tosses.
@@
In this paper, however, we allow the betting statistician to borrow interest-free before placing each bet or trade without being forced to quit the game in the possible event of indebtedness, and study if the emergence of evidence is changed in any way correspondingly. Further, we discuss the related scenario where the bets are underpriced such that arbitrage is possible without evidence.
Throughout this paper, we stick to the notation convention of using the letter $t$ to denote the discrete time index. The notation $\{x_{t}\}_{t\geqslant 1}$ refers to a discrete-time sequence $(x_{1},x_{2},\dots)$.
2 Problem Setup
2.1 The Classical Betting Game without Borrowing
@@
Let us first review the classical testing-by-betting setup pioneered by Shafer and Vovk [2005]. The setup involves the following game with two players, Casino [^1] This player is referred to as Forecaster in the original work, but in our work, Casino seems a better analogy. and Statistician. Casino tosses a coin repeatedly, realized as the random variables $X_{1},X_{2}\dots\in\{-1,1\}$, and Statistician does not believe that the coin is fair. For each $t\geqslant 1$, at time t1t-1, Casino offers a double-or-nothing bet to Statistician on the upcoming coin toss $X_{t}$: Statistician is allowed to put any fraction between 0 and 1 of their wealth on the table to bet on either heads ($X_{t}=1$) or tails ($X_{t}=-1$) of the outcome, and Casino pays Statistician double the amount on the table if Statistician made the correct guess. This corresponds to the following formal definition of the evolution of Statistician’s wealth process:
Definition 2.1 (Coin Betting).
Let $(\Omega,\mathcal{A},\mathbb{P})$ be a probability space with independent random variables $X_{t}=\Omega\to\{-1,1\}$ and filtration $\mathcal{F}_{t}=\sigma(X_{1},\dots,X_{t})$ for $t\geqslant 1$. Let $\{\lambda_{t}\}_{t\geqslant 1}$ be a [1,1][-1,1]-bounded stochastic process, predictable with respect to the filtration $\{\mathcal{F}_{t}\}_{t\geqslant 0}$. The wealth process $\{W_{t}\}_{t\geqslant 0}$ is defined as $W_{0}=1$ and
$W_{t}=W_{t-1}\cdot(1+\lambda_{t}X_{t}).$ (1)
The null hypothesis is that $\mathbb{P}(X_{t}=1)=1/2$ for all $t$. When this holds, we denote the probability by $\mathbb{P}_{\mathsf{M}}$ and the expected value by $\mathbb{E}_{\mathsf{M}}$.
Here, the sign of $\lambda_{t}$ corresponds to the side of Statistician’s bet: if $\lambda_{t}>0$, Statistician places $\lambda_{t}$ of their wealth to bet on heads $X_{t}=1$; if $\lambda_{t}<0$, $\lambda_{t}$ of their wealth on tails $X_{t}=-1$. We henceforth refer to the random variable $\lambda_{t}$ as the bet portion at time t1t-1, and the whole predictable sequence $\{\lambda_{t}\}_{t\geqslant 1}$ as the strategy. The null $\mathbb{P}_{\mathsf{M}}$ essentially states that the coin is fair with equal chance of heads and tails; or equivalently, the double-or-nothing odds that Casino offers are also fair. Since $\mathbb{E}_{\mathsf{M}}(X_{t})=0$, wealth $\{W_{t}\}$ is a nonnegative martingale under $\mathbb{P}_{\mathsf{M}}$ for any choice of the strategy $\{\lambda_{t}\}$: which is why we opt for the notation $\mathbb{P}_{\mathsf{M}}$.
The coin betting setup above is usually a primer to the more general setting of testing by a test supermartingale. We can replace the $1+\lambda_{t}X_{t}$ above by any nonnegative random variable with conditional expected value $\leqslant 1$. Formally, we have the following “NSM (nonnegative supermartingale) betting” setup of which Definition 2.1 is a very special case.
Definition 2.2 (NSM Betting).
Let $(\Omega,\mathcal{A},\mathbb{P})$ be a probability space with filtration $\{\mathcal{F}_{t}\}_{t\geqslant 0}$. Let $\{B_{t}\}_{t\geqslant 1}$ be a nonnegative process adapted to $\{\mathcal{F}_{t}\}_{t\geqslant 1}$. The wealth process $\{W_{t}\}_{t\geqslant 0}$ is defined as $W_{0}=1$ and
$W_{t}=W_{t-1}\cdot B_{t}.$ (2)
We consider two nested null hypotheses:
  • If $\mathbb{E}(B_{t}|\mathcal{F}_{t-1})\leqslant 1$ for all $t\geqslant 1$, we denote the probability by $\mathbb{P}_{\mathsf{S}}$ and the expected value by $\mathbb{E}_{\mathsf{S}}$. This is called the supermartingale null.
  • Further, if $\mathbb{E}(B_{t}|\mathcal{F}_{t-1})=1$ for all $t\geqslant 1$, we denote the probability by $\mathbb{P}_{\mathsf{M}}$ and the expected value by $\mathbb{E}_{\mathsf{M}}$. This is called the martingale null.
As a simple economic interpretation of Definition 2.2, $\{B_{t}\}$ is the random per-unit payoff at time $t$ of Statistician’s portfolio allocation at time t1t-1. Casino may offer various bets to Statistician, and Statistician may allocate their wealth $W_{t-1}$ arbitrarily between the bets offered (and cash). On aggregate, 1 unit of wealth invested in this portfolio at time t1t-1 becomes $B_{t}$ units at time $t$; and Statistician’s wealth consequently becomes $W_{t-1}\cdot B_{t}$ at time $t$. The null $\mathbb{P}_{\mathsf{M}}$ then states that the entire portfolio is always fairly priced, under which $\{W_{t}\}$ is a nonnegative martingale; The null $\mathbb{P}_{\mathsf{S}}$, on the other hand, states that the portfolio is at least fairly priced (with the possibility of being overpriced, to the disadvantage of Statistician), under which $\{W_{t}\}$ is a nonnegative supermartingale.
Since Definition 2.2 is a more general setup than Definition 2.1, our subsequent discussion as well as later development involving borrowing will build on Definition 2.2. We also keep the language of betting when we make this generalization: we shall talk of Statistician “betting $B_{t}$” in lieu of “betting on $X_{t}$ with bet portion $\lambda_{t}$” at time t1t-1. The quantity $B_{t}$ is referred to directly as the “bet at time t1t-1”. Sometimes, it will be helpful to think of the special coin-betting case $B_{t}=1+\lambda_{t}X_{t}$.
@@
In our setup, $\{W_{t}\}$ is a nonnegative supermartingale under $\mathbb{P}_{\mathsf{S}}$, and Ville’s inequality [Ville, 1939] then implies that, for all $x>1$,
$\mathbb{P}_{\mathsf{S}}\left(\sup_{t}W_{t}\geqslant x\right)\leqslant 1/x;$ (3)
@@
$\mathbb{P}_{\mathsf{S}}\left(W_{\tau}\geqslant x\right)\leqslant 1/x.$ (4)
Along with the optional stopping theorem for nonnegative supermartingales, which states that $\mathbb{E}_{\mathsf{S}}W_{\tau}\leqslant 1$ for any stopping time $\tau$, both inequalities above lend credibility to the interpretation that Statistician’s wealth $W_{t}$ can be seen as the evidence against $\mathbb{P}_{\mathsf{S}}$, the “subfairness” of the game to Statistician. Inequality (3) states that if the game were subfair, it would be unlikely for Statistician to ever be wealthy, and thus a large value of $W_{t}$ at any time means that the coin is unlikely to be fair. Inequality (4) states that Statistician may employ any stopping rule and a large final stopped wealth discredits $\mathbb{P}_{\mathsf{S}}$. Related to both is the concept of e-values, which are nonnegative random variables with expected values at most 1. In particular, $W_{\tau}$ is an e-value under $\mathbb{P}_{\mathsf{S}}$ for any stopping time $\tau$.
@@
2.2 Betting Game with Borrowing
We now suppose that before placing the bet at time t1t-1, Statistician borrows [^2] It does not matter from whom Statistician borrows the money. $\beta_{t}$ amount of money, increasing (or decreasing, if $\beta_{t}<0$) their current gross wealth $W_{t-1}$ to $W_{t-1}+\beta_{t}$, with which Statistician then bets $B_{t}$. Formally, we have the following generalization of Definition 2.2.
Definition 2.3 (Borrowed NSM Betting).
Let $(\Omega,\mathcal{A},\mathbb{P})$ be a probability space with filtration $\{\mathcal{F}_{t}\}_{t\geqslant 0}$. Let $\{B_{t}\}_{t\geqslant 1}$ be a nonnegative process adapted to $\{\mathcal{F}_{t}\}_{t\geqslant 1}$, and $\{\beta_{t}\}_{t\geqslant 1}$ a process predictable with respect to $\{\mathcal{F}_{t}\}_{t\geqslant 0}$ such that $\beta_{t}\geqslant-W_{t-1}$ for each $t$. The wealth process $\{W_{t}\}_{t\geqslant 0}$ is defined as $W_{0}=1$ and
$W_{t}=(W_{t-1}+\beta_{t})\cdot B_{t}.$ (5)
The supermartingale null $\mathbb{P}_{\mathsf{S}}$ under which $\mathbb{E}_{\mathsf{S}}(B_{t}|\mathcal{F}_{t-1})\leqslant 1$ for all $t\geqslant 1$, and the martingale null $\mathbb{P}_{\mathsf{M}}$ under which $\mathbb{E}_{\mathsf{M}}(B_{t}|\mathcal{F}_{t-1})=1$ for all $t\geqslant 1$ are as before.
The process $\{\beta_{t}\}$ is referred to as the borrowing strategy. In particular, we allow $\beta_{t}$ to be negative which could correspond to Statistician “paying back” earlier debts. To further clarify the set-up, we introduce the liabilities process denoting the total borrowed amount up to time $t$,
@@ 2.2 Betting Game with Borrowing
Recall that in the previous setting without borrowing in Section 2.1, the wealth process is a nonnegative martingale under $\mathbb{P}_{\mathsf{M}}$ and a nonnegative supermartingale under $\mathbb{P}_{\mathsf{S}}$. In our current case with borrowing, we have the following proposition on the more general relation between these three processes.
@@
The net wealth process $\{N_{t}\}$ is a martingale on $\{\mathcal{F}_{t}\}$ under $\mathbb{P}_{\mathsf{M}}$ and a supermartingale under $\mathbb{P}_{\mathsf{S}}$. Further, under $\mathbb{P}_{\mathsf{M}}$, $W_{t}=N_{t}+L_{t}$ is the Doob decomposition of the process $\{W_{t}\}$ into the martingale $\{N_{t}\}$ and the predictable process $\{L_{t}\}$.
@@
Note that
$\mathbb{E}(N_{t}|\mathcal{F}_{t-1})=(W_{t-1}+\beta_{t})\cdot\mathbb{E}(B_{t}|\mathcal{F}_{t-1})-L_{t}\begin{cases}=(W_{t-1}+\beta_{t})-L_{t}=N_{t-1},&\mathbb{E}=\mathbb{E}_{\mathsf{M}};\\ \leqslant(W_{t-1}+\beta_{t})-L_{t}=N_{t-1},&\mathbb{E}=\mathbb{E}_{\mathsf{S}}.\end{cases}$ (8)
therefore $\{N_{t}\}$ is a martingale under $\mathbb{P}_{\mathsf{M}}$ and supermartingale under $\mathbb{P}_{\mathsf{S}}$. It is easy to see that $\{L_{t}\}$ is predictable and $L_{0}=0$. Therefore, under $\mathbb{P}_{\mathsf{M}}$, $W_{t}=N_{t}+L_{t}$ is the Doob decomposition of $\{W_{t}\}$. ∎
3 Evidence in Borrowed Betting
The rest of the paper will be developed assuming the setup of Definition 2.3 until a further generalization is made in Section 9. We would like to study how Statistician, now holding both wealth $W_{t}$ and liabilities $L_{t}$, can quantify the evidence against $\mathbb{P}_{\mathsf{S}}$ or $\mathbb{P}_{\mathsf{M}}$, based on either $W_{t}$ or $N_{t}$ or both. To generalize our previous discussion on interpreting evidence via inequalities (3) and (4), let us introduce the following definitions.
@@ 3 Evidence in Borrowed Betting
Let $\tau$ be a stopping time on $\{\mathcal{F}_{t}\}$ and $a>0,b\geqslant 0$ be constants. We say that an $\mathcal{F}_{\tau}$-measurable random variable $E$ is an $(a,b)$-tail evidence against $\mathbb{P}$ at $\tau$ if the tail bound
$\mathbb{P}(E\geqslant x)\leqslant\frac{a}{x-b}$ (9)
@@
Random variables that are measurable with respect to the stopped $\sigma$-algebra $\mathcal{F}_{\tau}$ are often constructed by stopping at $\tau$ a process adapted to $\{\mathcal{F}_{t}\}$. This definition of tail evidence above follows the same reasoning that “if $\mathbb{P}$ holds, $E$ is unlikely to be large” as discussed in Section 2.1; the e-value $W_{\tau}$ back in (4), per Definition 3.1, is a $(1,0)$-tail evidence against $\mathbb{P}$ at $\tau$, and the constants $a,b$ in the definition are seen as “discount factors” of the evidence value $E$. More generally, if $a^{-1}(E-b)$ is an e-value, then $E$ is an $(a,b)$-tail evidence. Conversion from an $(a,b)$-tail evidence to a p-value for testing $\mathbb{P}$ can be done via $E\mapsto\frac{a}{(E-b)^{+}}$.
@@
$\mathbb{P}\left(\sup_{t}E_{t}\geqslant x\right)\leqslant\frac{a}{x-b},$ (10)
we say that it is a $(a,b)$-sequential tail evidence against $\mathbb{P}$.
For example, Ville’s inequality states that any nonnegative $\mathbb{P}$-supermartingale $\{M_{t}\}$ is an $(M_{0},0)$-sequential tail evidence against $\mathbb{P}$. The stopping time-dependent Definition 3.1 can be converted to the time-maximal Definition 3.2, if the same constant pair $(a,b)$ works for all bounded stopping times, as shown in the following proposition.
@@
Let $\{E_{t}\}$ be a process adapted to $\{\mathcal{F}_{t}\}$. It is an $(a,b)$-sequential tail evidence against $\mathbb{P}$ if and only if for any bounded stopping time $\tau$, $E_{\tau}$ is an $(a,b)$-tail evidence against $\mathbb{P}$ at $\tau$.
@@
First, suppose $E_{\tau}$ is an $(a,b)$-tail evidence at $\tau$ for any bounded stopping time $\tau$. Let $\sigma$ be the stopping time $\inf\{t:E_{t}\geqslant x\}$, where by convention $\inf\varnothing=\infty$. Then, since $E_{\sigma\wedge T}$ is an $(a,b)$-tail evidence against $\mathbb{P}$ at $\sigma\wedge T$,
$\mathbb{P}\left(\sup_{t\leqslant T}E_{t}\geqslant x\right)=\mathbb{P}(E_{\sigma\wedge T}\geqslant x)\leqslant\frac{a}{x-b}.$ (11)
@@
$\mathbb{P}\left(\sup_{t}E_{t}\geqslant x\right)=\mathbb{P}\left(\bigcup_{T=1}^{\infty}\left\{\sup_{t\leqslant T}E_{t}\geqslant x\right\}\right)\leqslant\frac{a}{x-b},$ (12)
@@
Next, suppose $\{E_{t}\}$ is an $(a,b)$-sequential tail evidence and let $\tau\leqslant T$ be a bounded stopping time. Then $\mathbb{P}(E_{\tau}\geqslant x)\leqslant\mathbb{P}(\sup_{t\leqslant T}E_{t}\geqslant x)\leqslant\frac{a}{x-b}$, concluding that $E_{\tau}$ is an $(a,b)$-tail evidence at $\tau$. ∎
@@
The net wealth process $\{N_{t}\}$ satisfies the supermartingale optional stopping theorem with respect to stopping times $\tau$ and 0 under $\mathbb{P}_{\mathsf{S}}$; that is, $\mathbb{E}_{\mathsf{S}}N_{\tau}\leqslant N_{0}=1$.